OpenStudy (anonymous):

IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. If a certain person has an IQ of 131, what percentage of the population has a lower IQ?

OpenStudy (kirbykirby):

Let \(X\) be the IQ score. If a person has \(X=131\), then the question asks what percentage (or what's the probability) that the population has a lower score. This is effectively asking for \(P(X<131)\). Since the scores are normally distributed (i.e. \(X\) follows a normal distribution), you can standardize \(X\) to obtain a standard normal distribution, and then use a standard normal table to find out the probability. Remember you standardize by doing: \[\frac{X-\mu}{\sigma} \], where \(\mu\) is the mean and \(\sigma\) is the standard deviation. So: \[P(X<131)=P\left( \frac{X-100}{15} < \frac{131-100}{15}\right)=P(Z<2.07) \]